# WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada

### v1

@@ 7 Conclusion
The fisrt limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate general-purpose models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
@@ Appendix A Data collection Queries
Table A.1 shows the full queries used to collect data from X.
@@
Table A.1: The full queries used for scraping data from X.
@@ Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic on our dataset, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^6] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
@@
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting a set number of topics to sort the data into, as at this stage, we were unsure of what sort of topics were contained in our data. BERTopic had sorted our data into over 100 fine-grained topics. This was still relatively large to manage, and it was difficult to determine a humanitarian category for each fine-grained topic. After some experimentation, we systematically decreased the total number of topics to determine which topics remained over time. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
Using BERTopic in the way described proved to be useful, as it gave us a good understanding from the beginning of what our future categories were to look like. Relying on this method alone was not enough, however, and having to look through the data manually at some point was inevitable. Many rounds of manual annotation ensued on a small portion of the data, after having it randomly mixed, in order to avoid constructing our categories on themes that may only exist at a certain time during the wildfire event. Through performing these steps, keeping in mind our discovered topics through BERTopic and humanitarian goals, new categories were added or dropped based on what was seen in the data. A final round of labelling then commenced, and the entire dataset was manually labeled.
@@ Appendix C Prompts for zero-shot classification
We show the full prompt used to evaluate VLMs in Table C.1.
Table C.1: System and User prompt used for gpt-4o-mini and LLaVa.
@@
"You are an assistant who is being given an image and text pair as a Twitter post which was created during a natural disaster event. Your task is to use information from both the text and the image to decide which option the post should be labeled as. You must pay close attention to each option when deciding which label to use."
@@ Appendix D Structures of custom classifiers
Appendix E Confusion matrix for all classifiers
For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figure E.1.
a) LLaVA, b) gpt-4o-mini, c) 2head-early, d) 3head-early, e) 2head-late, f) 3head-late
Figure E.1: Confusion matrices of classifiers and VLMs.

### v2

@@ 9 Acknowledgment
The first limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
@@
Finally, there is a broader risk of misapplication. Deploying these models in real-world disaster contexts without appropriate oversight, interpretability mechanisms, or human-in-the-loop verification may lead to misclassification, misinformation amplification, or missed signals, potentially affecting emergency coordination and public trust.
@@ Appendix A Data collection Queries
Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
@@
Table A.1: The queries used for collecting the data for WildFireCan-MMD.
@@ Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^11] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
@@
Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
@@ Appendix C Prompts for zero-shot classification
We show the full prompt used to evaluate the aforementioned VLMs in Table C.1. The following hyperparameters are used for all VLMs: temperature=0.1, max_new_tokens=1024.
Table C.1: System and User prompt used for VLMs.
@@
"You are an assistant who is being given an image and text pair as a Twitter post, which was created during a natural disaster event. Your task is to use information from both the text and the image to decide which option the post should be labeled as. You must pay close attention to each option when deciding which label to use."
@@ Appendix D Structures of custom classifiers
Appendix E Confusion matrices of classifiers
For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figures E.1 and E.2.
(a) 2-head classifier.
(b) 3-head classifier.
Figure E.1: Confusion matrices of custom classifiers.
(a) LLaVA.
(b) GPT-4.
(c) SmolVLM.
Figure E.2: Confusion matrices of VLMs.
Appendix F Queries Used for Additional Data Collection :
Below is the list of queries and keywords used to collect additional unlabeled data discussed in Section 6.
2018
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire)
Keyword Query: (alberta OR british columbia OR Prince George OR Grande Praire OR Waterton OR Bulkley Nechako OR Nadina Lake OR Kootenay OR Crowsnest Pass OR Medicine Lake OR Comstock Lake OR Tugwell Creek OR Sooke OR Nanaimo Lakes OR Tweedsmuir OR Johnny Creek OR Alkali Lake OR Lutz Creek OR Shovel Lake OR Nadina Lake OR Verdun Mountain OR Silver Lake OR Tommy Lakes OR Island Lake OR Chutanli Lake) (wildfire OR forest fire)
2019
Hashtag Query: (#ABWildfire OR #albertawildfire OR #ABFire)
Keyword Query: (alberta OR calgary OR edson OR Fort McMurray OR Grande Prairie OR High Level OR Lac La Biche OR Whitecourt OR Steen River OR Chuckegg Creek OR Peace River OR Slave Lake OR Wood Buffalo National Park) (wildfire OR forest fire)
2020
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire)
Keyword Query: (british columbia OR alberta OR sasksatchewan OR yukon OR northwest territories OR manitoba OR ontario OR quebec) (wildfire OR forest fire)
2021
Hashtag Query: (#MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #pafire OR #ontariofire OR #manitobafire OR #sasksatchewanfire)
Keyword Query: (manitoba OR ontario OR sasksatchewan OR british columbia) (wildfire OR forest fire)
2022
Hashtag Query: (#YTwildfire OR #yukonwildfire OR #YTfire OR #yukonforestfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #nwtforestfire)
Keyword Query: (#yzf OR #nwt OR #Yellowknife OR Yukon OR Northwest Territories OR Whitehorse OR Yellowknife OR Dawson City OR Great Slave Lake OR Norman Wells OR Inuvik OR Watson Lake OR Hay River OR Fort Smith OR Tuktoyaktuk OR Behchoko) (wildfire OR forest fire) (-alaska -Eielson -CityofNorthPole)
2023
Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
Keyword Query: (ontario OR quebec OR sasksatchewan OR british columbia OR manitoba OR northwest territories OR yukon OR alberta) (wildfire OR forest fire)
2024
Hashtag Query: (#JasperStrong OR #JasperWildfire OR #JasperAB OR #BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
Keyword Query: (canada OR ontario OR quebec OR sasksatchewan OR manitoba OR northwest territories OR yukon OR british columbia OR alberta OR jasper) (wildfire OR forest fire)
Note: Every query above includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
Appendix G Unlabeled dataset
A word cloud of the text data of the unlabeled dataset collected in Section 6, and also the frequency distribution of the posts by word count, are shown in Figures 1(a) and 1(b).
(a) Word cloud of the posts.
(b) Frquency distribution of the posts w.r.t their lengths.
Figure G.1: Word cloud of the textual part of the unlabeled tweets and the frequency distribution of them regarding their word counts.